OpenStudy (anonymous):

-112 squared question.. i used the factor tree and i am confused on how to pull the four out.. because i had a pair of 2 that i pulled out but thats the only pair? -_-

OpenStudy (saifoo.khan):

how about this? \[\sqrt{-1\times 2 \times 2 \times 2 \times 2 \times 7}\]

OpenStudy (anonymous):